Hi, could you please help me with this problem.
Solve:
2x+3y-z=5
2x+2y=7
8x-2y+3z=15
The answers are supposed to be: x = 0.786; y = 2.714 and z = 4.714.
i have gotten to every known number in the universe using all of my limited math skills, but not those three.
Please help.

Answer by Alan3354(69443)

Solve:
2x+3y-z=5
2x+2y=7
8x-2y+3z=15
The answers are supposed to be: x = 0.786; y = 2.714 and z = 4.714.
i have gotten to every known number in the universe using all of my limited math skills, but not those three.
Please help.
=======================
Think about this:
1/7 = 0.142857142857...